# hyperrefIgnoring empty anchor

### v1

@@ Abstract
We scrutinize the reported lensing anomaly of the CMB by considering several phenomenological modifications of the lensing consistency parameter, $A_{\rm L}$. Considering Planck spectra alone, we find statisically significant evidence for scale dependence (‘running’) of $A_{\rm L}$. We then demonstrate that the anomaly is entirely driven by Planck’s low multipoles, $\ell\leq 30$. When these data points are excluded a joint analysis with several other datasets clearly favors $\Lambda$CDM over the extended $\Lambda\rm CDM+A_{L}$ model. Not only that the lensing anomaly and low $\ell$ anomaly of the CMB go away in this case, but also the $S_{8}$ tension is ameliorated, and only the Hubble tension persists.
@@ I Introduction
The latter is a clear demonstration of the precision level achieved by present day cosmological probes. The simplicity of linear perturbations at the last scattering surface, as well as weak lensing of incoming CMB light by the intervening large scale structure, allow a precise calculation of the level at which the CMB spectrum is lensed given that we know the fundamental cosmological parameters. While it is indeed true that all we have at our disposal is the lensed sky, we can also capitalize on the fact that E-mode maps are significantly more sensitive to lensing than total intensity maps are. On the other hand, the latter are measured with a much higher S/N and so we gain a satisfactory inference of the cosmological parameters from intensity maps that allows us to forecast how the corresponding lensed E-mode maps should statistically look like. With the exquisite measurements provided by the Planck satellite, it was deduced that the observed lensing signal is $\gtrsim 2\sigma$ larger than expected assuming the flat $\Lambda$CDM model. As we show below, this anomaly is chiefly contributed by the lowest $l\lesssim 30$ multipoles, where – perhaps significantly – most of the evidence for nonvanishing spatial curvature evidence comes from, as well as the infamous ‘$-\ell$ anomaly’ [12, 13, 14].
@@
Every cosmological observable, e.g. CMB power spectra, galaxy correlation, shear maps, etc., has its own degeneracies between the various cosmological parameters. All the anomalies are parameterized by cosmological parameters and it then follows that cosmological anomalies are generally correlated. This well-known fact is nicely demonstrated in Figure 1. Thus, removing some anomaly by reducing systemtaic errors, or invoking New Physics, can considerably influence all other anomalies for better or for worse. In this paper we adopt a phenomenology driven approach to the lensing anomaly, using several techniques, such as allowing for scale dependence of the lensing anomaly, binning and removing various subsets of the data, and combining various data sets. In these analyses, we attempted at simultaneously inferring the cosmological parameters of $\Lambda$CDM, and the extended $\Lambda$CDM model, that includes the lensing amplitude, $A_{\rm L}$, which is denoted as $\Lambda$CDM$+A_{\rm L}$ throughout. In addition, we analyze an extended cosmological model that includes scale dependence of the lensing with either a weak logarithmic or power-law dependencies on the angular scale. We also analyze an extension considering a new consistency parameter for lensing of EE modes independently, denoted as $A_{EE}$, such that the extended model has 7 parameters $\Lambda$CDM$+A_{\rm EE}$.
@@
  • We considered various combinations of CMB data sets [59, 60, 61, 62] and in one case we also included low redshift data [44, 63, 64, 65, 66]. In all data set combinations considered, we removed the low multipoles $\ell\leq 30$ Planck data. As a result, both the lensing anomaly and the low-ell anomaly are gone, such that $A_{L}=1$ is consistent within $1\sigma$, in accord with other analyses [62, 67, 68]. By adopting a certain model selection criterion, as is detailed in the next section, we find that the data (with $-\ell$ data excluded) favors the $\Lambda$CDM model over its extension $\Lambda$CDM+$A_{L}$. Removing Planck data completely also reduces the $S_{8}$ tension, and only the Hubble tension persists at its reported $\gtrsim 4\sigma$ level.
The paper is organized as follows. We begin by defining the lensing consistency amplitude $A_{L}$ and discuss the current status of the anomaly in section II. The different data sets and methodology are specified and described in section III. We describe the results of splitting and binning the multipoles in section IV and V, as well as demonstrating that the anomaly is driven by the low multipoles. The effect of a new consistency parameter $A_{EE}$ is discussed in section VI. In section VII we analyze the possibility of a scale dependent consistency parameter, and its statistical significance. Finally, we remove the ‘culprit’ multipoles, and include the additional data sets, in section VIII. We demonstrate that the the lensing anomaly goes away in the revised analysis, i.e. our global parameter estimation is consistent with $A_{L}=1$. We then conclude in section IX.
@@ II The lensing anomaly
where $\hat{\textbf{n}}$ is radially inward photon direction, $\Psi$ is the well-known Weyl potential of the scalar perturbation theory, $\chi_{*}$ is the conformal distance of decoupling, and $\eta_{0}-\chi$ is the conformal time at which the photon was at position $\chi.\hat{\textbf{n}}$. Using the Fourier transform of Weyl potential:
@@
$\tilde{C}_{l}^{\Theta}\approx\left(1-l^{2}R^{\psi}\right)C_{l}^{\Theta}+\int\frac{\mathrm{d}^{2}\mathbf{l}^{\prime}}{(2\pi)^{2}}\left[\mathbf{l}^{\prime}\cdot\left(\mathbf{l}-\mathbf{l}^{\prime}\right)\right]^{2}C_{\left|\mathbf{l}-\mathbf{l}^{\prime}\right|}^{\psi}C_{l^{\prime}}^{\Theta},$ (6)
where $\tilde{C}_{l}^{\Theta}$ is the lensed- and $C_{l}^{\Theta}$ is the unlensed power spectrum and $R^{\psi}=\frac{1}{4\pi}\int\frac{dl}{l}l^{4}C_{l}^{\psi}$. Weak lensing is phenomenologically decoupled from the primary anisotropies by introducing a parameter $A_{L}$, [73] i.e.,
@@ III Data-Sets and Methodology
    We make use of the measurements from the nine-year WMAP temperature, polarization, and lensing maps (referred to as WMAP) [59] excluding the TE data below $\ell=24$ to utilize a Gaussian prior on $\tau$. We incorporate the latest datasets of ACT DR4 TTTEEE [61, 62] with the combination of ACT DR6 Lensing (the combined dataset, DR4+DR6, is referred to as ACT) [91, 67] and SPT-3G 2018 TT/TE/EE (referred to as SPT)[60].
@@
    Large Scale Structure data provides very decisive observational constraints on the growth of structure in our Universe. Baryon Acoustic Oscillation (BAO) and Redshift space distortions (RSD) measurements from the SDSS-IV eBOSS survey include isotropic and anisotropic distance, expansion rate measurements, and measurements of $f\sigma 8$. We use the Seventh Data Release of the SDSS Main Galaxy Sample [44], final clustering measurements from the eBOSS survey [63], and the SDSS’s Twelfth and Sixteenth Data Release [64, 65]. These datasets include measurements of Luminous Red Galaxies (LRG), Emission Line Galaxies (ELG), Quasars (QSO), the Lyman-alpha forest auto-correlation (lyauto), and the Lyman-alpha Quasar cross-correlation (lyxqso). Moreover, we include the PantheonPlus dataset [66] containing 1701 light curves for 1550 spectroscopically confirmed Type Ia supernovae (SNeIa) in the redshift range $2.260.001<z<$ 2.26.
@@
The prior ranges we adopted for the seven cosmological parameters are summarized in Table 1. Specifically, we set a wide flat prior for $A_{L}\in[0.1,5]$. Priors related to further extensions of the cosmological model will be specified in their relevant section. To make a quantitative measure of how well a given model performs with a given dataset we make use of the deviance information criterion, DIC, defined as:
@@
where $\theta$ is the model parameter vector, and the bar represents averages over the posterior distribution $\mathcal{P}(\theta)$. For any two models M1 and M2 we use $\Delta\rm DIC=DIC_{M1}-DIC_{M2}$ to quantify the evidence of the dataset favoring a model. Popularly, $-2\leq\Delta\rm DIC\leq 0$ is considered as weak preference of M1 over M2, $-6\leq\Delta\rm DIC<-2$ indicates that that data moderately favors M1 over M2, $-10\leq\Delta\rm DIC<-6$ is usually taken as a strong evidence that M1 outperforms M2, and $\Delta\rm DIC<-10$ is considered as a very strong evidence for the preference of M1 over M2 [95, 81].
@@ IV Multipole Splitting
Apart from one outlier, we find that the inferred values of the lensing parameter are consistent with P18 full data results, as evident from the top figure of the left panel of Fig. 2. We emphasize that the variation in the inferred value of the lensing amplitude can affect other cosmological parameters, specifically $H_{0}$ and $S_{8}$ as shown in middle and bottom sub-panels of the left panel in Fig. 2 respectively. We note a marginal increase in derived values of $H_{0}$ but overall consistent with the complete P18 data using $\Lambda$CDM and its extension with $A_{L}$. However, this consistency of the former with the latter is due to increased statistical uncertainties. Finally, $S_{8}$ has no significant influence on the variation of lensing amplitude over multipole bins $\in[2,l_{\text{max}}]$.
@@ V Binned $A_{\text{L}}$
In the previous section, we examined the impact of low and high multipoles on the inferred $A_{L}$ by excluding data at the high- or low-end of multipoles. However, this splitting of the CMB data comes at a cost of increased statistical uncertainties on the inferred cosmological parameters. In this section, we introduce a new approach to investigate the impact of low $\ell$ (or high $\ell$) while utilizing the complete Planck dataset. The gravitational lensing parameter $A_{\text{L}}$ was originally introduced as a constant parameter effective at all $\ ell$s. In this section, we study the impact of lensing at different multipole bins. To allow for $\ell$-dependent lensing anomaly, we define the lensing parameter as follows:
@@
In Equation 9, $A_{\text{L}}$ is a free parameter when $\ell$ lies within the range $[\ell_{\text{min}},\ell_{\text{max}}]$, and it is set to unity otherwise. We modify CAMB [93] to effect this departure from the standard treatment. The underlying rational of such a construction is two-fold. First, we are looking for evidence of a deviation from the expected gravitational lensing at a particular $\ell$ range. Second, such a parametrization of $A_{\text{L}}$ could potentially offer a resolution of several pronounced cosmological tensions, including Hubble, Large Scale Structure, and $A_{\text{L}}$ tensions simply because we suspect that a few of these tensions are correlated and so a more accurate modelling of the lensing anomaly might shed a brighter light on the origin of other – correlated – anomalies. We analyze the model in two ways, depending on the choice of $\ell_{max}$ and $\ell_{min}$, described as follows:
  • We fix the $\ell_{min}=2$ and increase $\ell_{max}$ from 100 to 2500 with $\Delta\ell=100$. We refer to this analysis approach as Type I.
  • We start with $\ell_{min}=2$ and $\ell_{max}=30$ and increase both $\ell_{min}$ and $\ell_{max}$ with $\Delta\ell=100$ up to $\ell_{max}=2500$. We refer to this approach as Type II.
@@
Fig. 3 shows the 68% CL inference for $A_{\text{L}},H_{0}$ and $S_{8}$ parameters for different values of multipoles in Type I. We compare the constraints on $H_{0}$, $S_{8}$, and $A_{\text{L}}$ using primary P18 measurements, both including and excluding the lensing likelihood. Inclusion of lensing likelihood with primary P18 anisotropies, the inferred values of $A_{\text{L}}$ remain unchanged from different $\ell$ ranges. However, when the lensing likelihood is excluded, the lensing amplitude increases across the entire $\ell$ range. This increase in the inferred value of $A_{\text{L}}$ is most significant for $\ell\in[2,100]$. Other cosmological parameters $H_{0}$ and $S_{8}$ are consistent with Planck values - with and without the lensing likelihood. In particular, we note that now the value of the Hubble parameter deviates slightly from its prediction by the standard model using Planck data, while $\sigma_{8}$ remains consistent with the standard model inference.
@@ VI Separating the polarized and unpolarized parts of the lensing anomaly
As E-modes seem to have a significant impact on the lensing amplitude, see also [97], we introduce a separate consistency parameter for E-modes. This is motivated also by the fact that the underlying source for the anomalous level of observed lensing is not yet known and so there is no justification to apply the same lensing anomaly parameter to both temperature anisotropy and E-mode polarization other than simplicity. In fact, it is not hard to envisage a scenario in which temperature anisotropy and E-mode polarization might require two different anomaly parameter. For example, the relatively E-mode signal is much more prone to all types of systematics, foregrounds, etc. than the temperature anisotropy does, and certain optical imperfections of the telescope could easily leak (the much larger) unpolarized intensity to polarization maps. On the other hand, polarized signal is especially sensitive to lensing, and so the lensing anomaly could in principle result from contaminated E-mode signal. By applying the same anomaly parameter, $A_{L}$, as is conventionally done, to both temperature anisotropy and polarization the inferred lensing anomaly is potentially systematically biased towards lower values. By separating the two contributions to this anomaly we relax the linkage between T and E, and in principle allow the contribution of temperature anisotropy to the lensing anomaly to be diminishing while at the same time increase the contribution of the E-mode polarization to the lensing anomaly. Since we have no a priori prejudice as to the likely source of this anomaly we would ideally allow for a parameterization that is as bias-free as possible.
We introduce a phenomenological parameter which re-scales the EE and TE CMB power spectrum by a factor of $\text{A}_{\text{EE}}$. To do so, we multiply the theoretical prediction of E modes power spectrum in CAMB by $\text{A}_{\text{EE}}$. The modification has been implemented such that the EE and TE spectra are treated differently. Specifically, the $C_{l}^{EE}$ spectrum is scaled by $A_{EE}$, while the $C_{l}^{TE}$ spectrum is scaled by $\sqrt{A_{EE}}$. This approach ensures that the spectra are adjusted correctly, with $C_{l}^{EE}$ transforming as $C_{l}^{EE}\rightarrow A_{EE}C_{l}^{EE}$ and $C_{l}^{TE}$ as $C_{l}^{TE}\rightarrow\sqrt{A_{EE}}C_{l}^{TE}$. Fig. 5 shows the dependence of CMB’s EE and TE power spectra on the new consistency parameters. Other cross-correlated CMB anisotropies remain unaffected by the new consistency parameter $A_{\text{EE}}$. We shall be interested in inferring of the new parameter $\text{A}_{\text{EE}}$ using CMB experiments including P18, WMAP, ACT and SPT-3G.
@@
We conclude our analysis of P18 data for the $\Lambda$CDM + $\text{A}_{\text{EE}}$ case by only considering $-\ell$ EE modes. Interestingly, we obtain $\text{A}_{\text{EE}}=1.59^{+0.51}_{-0.65}$ parameter at 68% CL showing the agreement with its canonical value. We further point out that the consideration of only $\ell$ EE spectrum gives slightly better $\Delta DIC=-2.19$ which marginally better as compared to other analyses presented in Table 3. However, due to significantly less data provided by low $\ell$ EE spectrum, a few of the cosmological parameters remain unconstrained. In left panel of Fig. 6, we show the posterior distributions of cosmological parameters $H_{0}$, $S_{8}$ and $A_{EE}$ for $\Lambda$CDM and $\Lambda$CDM + $A_{\rm EE}$ model using all P18 data. As shown in Table 3, the optical depth $\tau$ is influenced by new $A_{EE}$ parameter. Therefore, we also include the contour plots of $\tau$ and $A_{EE}$ in right panel of Fig. 6. We show our results for Planck $-\ell$ EE and $-\ell$ TTTEEE + $-\ell$ EE data in green and blue respectively.

### v2

@@ Abstract
We scrutinize the reported lensing anomaly of the CMB by considering several phenomenological modifications of the lensing consistency parameter, $A_{\rm L}$. Considering Planck spectra alone, we find statistically significant evidence for scale dependence (‘running’) of $A_{\rm L}$. We then demonstrate that the anomaly is entirely driven by Planck’s low multipoles, $\ell\leq 30$. When these data points are excluded, a joint analysis with several other datasets clearly favors $\Lambda$CDM over the extended $\Lambda\rm CDM+A_{L}$ model. Not only that the lensing anomaly and low $\ell$ anomaly of the CMB go away in this case, but also the $S_{8}$ tension is ameliorated, and only the Hubble tension persists.
@@ I Introduction
The latter is a clear demonstration of the precision level achieved by present day cosmological probes. The simplicity of linear perturbations at the last scattering surface, as well as weak lensing of incoming CMB light by the intervening large scale structure, allow a precise calculation of the level at which the CMB spectrum is lensed given that we know the fundamental cosmological parameters. While it is indeed true that all we have at our disposal is the lensed sky, we can also capitalize on the fact that E-mode maps are significantly more sensitive to lensing than total intensity maps are. On the other hand, the latter are measured with a much higher S/N and so we gain a satisfactory inference of the cosmological parameters from intensity maps that allows us to forecast how the corresponding lensed E-mode maps should statistically look like. With the exquisite measurements provided by the Planck satellite, it was deduced that the observed lensing signal is $\gtrsim 2\sigma$ larger than expected assuming the flat $\Lambda$CDM model. As we show below, this anomaly is chiefly contributed by the lowest $\ell\lesssim 30$ multipoles, where – perhaps significantly – most of the evidence for nonvanishing spatial curvature evidence comes from, as well as the infamous ‘$-\ell$ anomaly’ [12, 13, 14].
@@
Every cosmological observable, e.g. CMB power spectra, galaxy correlation, shear maps, etc., has its own degeneracies between the various cosmological parameters. All the anomalies are parameterized by cosmological parameters and it then follows that cosmological anomalies are generally correlated. This well-known fact is nicely demonstrated in Fig. 1. Thus, removing some anomaly by reducing systemtaic errors, or invoking New Physics, can considerably influence all other anomalies for better or for worse. In this paper we adopt a phenomenology driven approach to the lensing anomaly, using several techniques, such as allowing for scale dependence of the lensing anomaly, binning and removing various subsets of the data, and combining various data sets. In these analyses, we attempted at simultaneously inferring the cosmological parameters of $\Lambda$CDM, and the extended $\Lambda$CDM model, that includes the lensing amplitude, $A_{\rm L}$, which is denoted as $\Lambda$CDM$+A_{\rm L}$ throughout. In addition, we analyze an extended cosmological model that includes scale dependence of the lensing with either a weak logarithmic or power-law dependencies on the angular scale. We also analyze an extension considering a new consistency parameter for lensing of EE modes independently, denoted as $A_{EE}$, such that the extended model has 7 parameters $\Lambda$CDM$+A_{\rm EE}$.
@@
  • We considered various combinations of CMB data sets [59, 60, 61, 62] and in one case we also included low redshift data [44, 63, 64, 65, 66]. In all data set combinations considered, we removed the low multipoles $\ell\leq 30$ Planck data. As a result, both the lensing anomaly and the $-\ell$ anomaly are gone, such that $A_{L}=1$ is consistent within $1\sigma$, in accord with other analyses [62, 67, 68]. By adopting a certain model selection criterion, as is detailed in the next section, we find that the data (with $-\ell$ data excluded) favors the $\Lambda$CDM model over its extension $\Lambda$CDM+$A_{L}$. Removing Planck data completely (while retaining other CMB data) also reduces the $S_{8}$ tension, and only the Hubble tension persists at its reported $\gtrsim 4\sigma$ level.
The paper is organized as follows. We begin by defining the lensing consistency amplitude $A_{L}$ and discuss the current status of the anomaly in section II. The different data sets and methodology are specified and described in section III. We describe the results of splitting and binning the multipoles in section IV and V, as well as demonstrating that the anomaly is driven by the low multipoles. The effect of a new consistency parameter $A_{EE}$ is discussed in section VI. In section VII we analyze the possibility of a scale dependent consistency parameter, and its statistical significance. Finally, we remove the ‘culprit’ multipoles, and include the additional data sets, in section VIII. We demonstrate that the lensing anomaly goes away in the revised analysis, i.e. our global parameter estimation is consistent with $A_{L}=1$. We then conclude in section IX.
@@ II The lensing anomaly
where $\hat{\textbf{n}}$ is radially inward photon direction, $\Psi$ is the well-known Weyl potential of the scalar perturbation theory, $\chi_{*}$ is the conformal distance of decoupling, and $\eta_{0}-\chi$ is the conformal time at which the photon was at position $\chi.\hat{\textbf{n}}$. Using the Fourier transform of the Weyl potential:
@@
$\tilde{C}_{\ell}^{\Theta}\approx\left(1-\ell^{2}R^{\psi}\right)C_{\ell}^{\Theta}+\int\frac{\mathrm{d}^{2}\mathbf{\ell}^{\prime}}{(2\pi)^{2}}\left[\mathbf{\ell}^{\prime}\cdot\left(\mathbf{\ell}-\mathbf{\ell}^{\prime}\right)\right]^{2}C_{\left|\mathbf{\ell}-\mathbf{\ell}^{\prime}\right|}^{\psi}C_{\ell^{\prime}}^{\Theta},$ (6)
where $\tilde{C}_{\ell}^{\Theta}$ is the lensed- and $C_{\ell}^{\Theta}$ is the unlensed power spectrum, $\Theta(={\delta T}/T)$ being the temperature field of the CMB, and $R^{\psi}=\frac{1}{4\pi}\int\frac{d\ell}{\ell}\ell^{4}C_{\ell}^{\psi}$. Lensing of CMB photons also affects the polarization modes; for details, see reference [72]. Weak lensing is phenomenologically decoupled from the primary anisotropies by introducing a parameter $A_{L}$, [73] i.e.,
@@ III Data-Sets and Methodology
    We make use of the measurements from the nine-year WMAP temperature, polarization, and lensing maps (referred to as WMAP) [59] excluding the TE data below $\ell=24$ utilizing a Gaussian prior on $\tau$. We incorporate the latest datasets of ACT DR4 TTTEEE [61, 62] with the combination of ACT DR6 Lensing (the combined dataset, DR4+DR6, is referred to as ACT) [91, 67] and SPT-3G 2018 TT/TE/EE (referred to as SPT)[60].
@@
    Large Scale Structure data provides very decisive observational constraints on the growth of structure in our Universe. Baryon Acoustic Oscillation (BAO) and Redshift space distortions (RSD) measurements from the SDSS-IV eBOSS survey include isotropic and anisotropic distance, expansion rate measurements, and measurements of $f\sigma_{8}$. We use the Seventh Data Release of the SDSS Main Galaxy Sample [44], final clustering measurements from the eBOSS survey [63], and the SDSS’s Twelfth and Sixteenth Data Release [64, 65]. These datasets include measurements of Luminous Red Galaxies (LRG), Emission Line Galaxies (ELG), Quasars (QSO), the Lyman-alpha forest auto-correlation (lyauto), and the Lyman-alpha Quasar cross-correlation (lyxqso). Moreover, we include the PantheonPlus dataset [66] containing 1701 light curves for 1550 spectroscopically confirmed Type Ia supernovae (SNeIa) in the redshift range $2.260.001<z<$ 2.26.
@@
The prior ranges we adopted for the seven cosmological parameters are summarized in Table 1. Specifically, we set a wide flat prior for $A_{L}\in[0.1,5]$. Priors related to further extensions of the cosmological model will be specified in their relevant section. To make a quantitative measure of how well a given model performs with a given dataset we make use of the deviance information criterion, DIC, defined as [95, 96]:
@@
where $\theta$ is the model parameter vector, and the bar represents averages over the posterior distribution $\mathcal{P}(\theta)$. For any two models M1 and M2 we use $\Delta\rm DIC=DIC_{M1}-DIC_{M2}$ to quantify the evidence of the dataset favoring a model. Popularly, $-2\leq\Delta\rm DIC\leq 0$ is considered as weak preference of M1 over M2, $-6\leq\Delta\rm DIC<-2$ indicates that that data moderately favors M1 over M2, $-10\leq\Delta\rm DIC<-6$ is usually taken as a strong evidence that M1 outperforms M2, and $\Delta\rm DIC<-10$ is considered as a very strong evidence for the preference of M1 over M2 [97, 98, 99, 81].
@@ IV Multipole Splitting
Apart from one outlier, we find that the inferred values of the lensing parameter are consistent with P18 full data results, as evident from the top figure of the left panel of Fig. 2. We emphasize that the variation in the inferred value of the lensing amplitude can affect other cosmological parameters, specifically $H_{0}$ and $S_{8}$ as shown in middle and bottom sub-panels of the left panel in Fig. 2 respectively. We note a marginal increase in derived values of $H_{0}$ but overall consistent with the complete P18 data using $\Lambda$CDM and its extension with $A_{L}$. However, this consistency of the former with the latter is due to increased statistical uncertainties. Finally, $S_{8}$ has no significant influence on the variation of lensing amplitude over multipole bins $\in[2,\ell_{\text{max}}]$.
@@ V Binned $A_{\text{L}}$
In the previous section, we examined the impact of low and high multipoles on the inferred $A_{L}$ by excluding data at the high- or low-end of multipoles. However, this splitting of the CMB data comes at a cost of increased statistical uncertainties on the inferred cosmological parameters. In this section, we introduce a new approach to investigate the impact of low $\ell$ (or high $\ell$) while utilizing the complete Planck dataset. The gravitational lensing parameter $A_{\text{L}}$ was originally introduced as a constant parameter effective at all $\ell$s. Instead, we study the impact of lensing at different multipole bins. To allow for $\ell$-dependent lensing anomaly, we define the lensing parameter as follows:
@@
In Eq. 9, $A_{\text{L}}$ is a free parameter when $\ell$ lies within the range $[\ell_{\text{min}},\ell_{\text{max}}]$, and it is set to unity otherwise. We modify CAMB [93] to effect this departure from the standard treatment. The underlying rational of such a construction is two-fold. First, we are looking for evidence of a deviation from the expected gravitational lensing at a particular $\ell$ range. Second, such a parametrization of $A_{\text{L}}$ could potentially offer a resolution of several pronounced cosmological tensions, including Hubble, Large Scale Structure, and $A_{\text{L}}$ tensions simply because we suspect that a few of these tensions are correlated and so a more accurate modelling of the lensing anomaly might shed a brighter light on the origin of other – correlated – anomalies. We analyze the model in two ways, depending on the choice of $\ell_{\max}$ and $\ell_{\min}$, described as follows:
  • We fix the $\ell_{\min}=2$ and increase $\ell_{\max}$ from 100 to 2500 with $\Delta\ell=100$. We refer to this analysis approach as Type I.
  • We start with $\ell_{\min}=2$ and $\ell_{\max}=30$ and increase both $\ell_{\min}$ and $\ell_{\max}$ with $\Delta\ell=100$ up to $\ell_{\max}=2500$. We refer to this approach as Type II.
@@
Fig. 3 shows the 68% CL inference for $A_{\text{L}},H_{0}$ and $S_{8}$ parameters for different values of multipoles in Type I. We compare the constraints on $H_{0}$, $S_{8}$, and $A_{\text{L}}$ using primary P18 measurements, both including and excluding the lensing likelihood. Inclusion of lensing likelihood with primary P18 anisotropies, the inferred values of $A_{\text{L}}$ remain unchanged from different $\ell$ ranges. However, when the lensing likelihood is excluded, the lensing amplitude increases across the entire $\ell$ range. This increase in the inferred value of $A_{\text{L}}$ is most significant for $\ell\in[2,100]$. Other cosmological parameters like $H_{0}$ and $S_{8}$ are consistent with Planck values - with and without the lensing likelihood. In particular, we note that now the value of the Hubble parameter deviates slightly from its prediction by the standard model using Planck data, while $\sigma_{8}$ remains consistent with the standard model inference.
@@ VI Separating the polarized and unpolarized parts of the lensing anomaly
As E-modes seem to have a significant impact on the lensing amplitude, see also [101], we introduce a separate consistency parameter for E-modes. This is motivated also by the fact that the underlying source for the anomalous level of observed lensing is not yet known and so there is no justification to apply the same lensing anomaly parameter to both temperature anisotropy and E-mode polarization other than simplicity. In fact, it is not hard to envisage a scenario in which temperature anisotropy and E-mode polarization might require two different consistency parameters. For example, the E-mode signal is much more prone to all types of systematics, foregrounds, etc. than the temperature anisotropy does, and certain optical imperfections of the telescope could easily leak (the much larger) unpolarized intensity to polarization maps. On the other hand, polarized signal is especially sensitive to lensing, and so the lensing anomaly could in principle result from contaminated E-mode signal. By applying the same anomaly parameter, $A_{L}$, as is conventionally done, to both temperature anisotropy and polarization the inferred lensing anomaly is potentially systematically biased towards lower values. By separating the two contributions to this anomaly we relax the linkage between T and E, and in principle allow the contribution of temperature anisotropy to the lensing anomaly to be diminishing while at the same time increase the contribution of the E-mode polarization to the lensing anomaly. Since we have no a priori prejudice as to the likely source of this anomaly we would ideally allow for a parameterization that is as bias-free as possible.
We introduce a phenomenological parameter which re-scales the EE and TE CMB power spectrum by a factor of $\text{A}_{\text{EE}}$. To do so, we multiply the theoretical prediction of E modes power spectrum in CAMB by $\text{A}_{\text{EE}}$. The modification has been implemented such that the EE and TE spectra are treated differently. Specifically, the $C_{\ell}^{EE}$ spectrum is scaled by $A_{EE}$, while the $C_{\ell}^{TE}$ spectrum is scaled by $\sqrt{A_{EE}}$. This approach ensures that the spectra are adjusted correctly, with $C_{\ell}^{EE}$ transforming as $C_{\ell}^{EE}\rightarrow A_{EE}C_{\ell}^{EE}$ and $C_{\ell}^{TE}$ as $C_{\ell}^{TE}\rightarrow\sqrt{A_{EE}}C_{\ell}^{TE}$. Fig. 5 shows the dependence of CMB’s EE and TE power spectra on the new consistency parameters. Other cross-correlated CMB anisotropies remain unaffected by the new consistency parameter $A_{\text{EE}}$. We shall be interested in inferring of the new parameter $\text{A}_{\text{EE}}$ using CMB experiments including P18, WMAP, ACT and SPT-3G.
@@
We conclude our analysis of P18 data for the $\Lambda$CDM + $\text{A}_{\text{EE}}$ case by only considering $-\ell$ EE modes. Interestingly, we obtain $\text{A}_{\text{EE}}=1.59^{+0.51}_{-0.65}$ parameter at 68% CL showing the agreement with its canonical value. We further point out that the consideration of only $\ell$ EE spectrum gives slightly better $\Delta\rm DIC=-2.19$ which marginally better as compared to other analyses presented in Table 3. However, due to significantly less data provided by low $\ell$ EE spectrum, a few of the cosmological parameters remain unconstrained. In left panel of Fig. 6, we show the posterior distributions of cosmological parameters $H_{0}$, $S_{8}$ and $A_{EE}$ for $\Lambda$CDM and $\Lambda$CDM + $A_{\rm EE}$ model using all P18 data. As shown in Table 3, the optical depth $\tau$ is influenced by new $A_{EE}$ parameter. Therefore, we also include the contour plots of $\tau$ and $A_{EE}$ in right panel of Fig. 6. We show our results for Planck $-\ell$ EE and $-\ell$ TTTEEE + $-\ell$ EE data in green and blue respectively.